OpenStudy (anonymous):

can someone explain this is my homework 17 = –13 – 8x

OpenStudy (anonymous):

An equation remains true if you do the same thing to both sides. Start by getting rid of the -13. To do this, add 13 to BOTH SIDES of the equation: now you have: 17 + 13 = -13 - 8x + 13 On the left side, 17 + 13 = 30. On the right side, the -13 and +13 cancel, leaving just the -8x beind. 30 = -8x Now, we want to find a number that equals thirty when you multiply by -8. To do this, just do 30/-8 = -15/4. So x = -15/4

OpenStudy (anonymous):